OpenStudy (liljus13):

Variable y varies directly with x2, and y = 48 when x = 4. Which graph represents the quadratic variation?

OpenStudy (liljus13):

help

OpenStudy (mathmale):

Please express that square as x^2 (rather than x2). If y varies directly with x^2, and if y=48 when x=4, graph the quadratic equation (since you haven't shared the ones that came with this problem): Let k be the constant of (direct) proportionality. Then y = k x^2. substitute y=48 and x=4 and solve for k. Then write the above quadratic equation again, using your value of k. Graph y=x^2. How would the graph change if you then wrote y=2x^2, or y=3x^2?